OpenStudy (anonymous):

A cylinder with radius r and height 2r +4 contains a cube with edge length r times square root of 2. What fraction of the cylinder's volume is taken up by the cube. Write in simplified form.

OpenStudy (anonymous):

V_cylinder = (Pi r^2) (2r+4) V_cube = (sqrt(2) r)^3 {I assume that's what you meant.} Fraction = V_cyl / V_cube = (2Pi r^3 + 4Pi r^2) / (2sqrt(2) r^3) = (2Pi r + 4Pi) / 2sqrt(2) r = Pi (r+2) / sqrt(2) r = Pi sqrt(2) (r+2) / 2r Ed: Oops - I inverted this... the correct answer is sqrt(2) r / Pi (r+2)

OpenStudy (anonymous):

thank you!!! You're the best!

OpenStudy (anonymous):

Np, No you are :D.

OpenStudy (anonymous):

Glad i could help :)!